Question 3, chap 122, sect 1. part of 10 points The efficiency of an 880 MW nuclear power plant is 16.2% If a river of flow rate 2.69 x 106 kg/s were used to transport the excess thermal energy away, what would be the average temperature increase of the river? Answer in units of C.
Answer #1

The excess thermal energy transported per second by the river is

Pexcess=Pinput(1-n)=(Poutput/n)(1-n)

Pexcess=(880/0.162)(1-0.162)=4552.1 MW

since

Pexcess=c(dm/dt)*dT

4552.1*106=(4184)(2.69*106)*dT

dT=0.404oC